Referências

[1] Zvi Artstein. Examples of stabilization with hybrid feedback. In R Alur, TA Henzinger, and ED Sontag, editors, Hybrid System III, volume 1066 of Lect. Notes Comp. Sci., chapter 13, pages 173–185. Springer Verlag, Berlin, 1996.

[2] K. J. Åström and B. Wittenmark. Adaptive Control. Addison-Wesley, 1989.

[3] S Bennett. A history of control engineering, 1800-1930. Peter Peregrinus / IEE, London, 1979.

[4] S Bennett. A history of control engineering, 1930-1955. Peter Peregrinus / IEE, London, 1993.

[5] Michael S Branicky. Studies in Hybrid Systems: Modelling, Analysis, and Control. PhD thesis, Massachussetts Institute of Technology, June 1995.

[6] Wen-Chung Chang, AS Morse, and Gregory D Hager. A calibrationfree, self-adjusting stereo visual control system. In 13th World Congress, pages 343–348, San Francisco, July 1996. IFAC.

[7] Simon Haykin. Adaptive filter theory. Prentice Hall, Englewood Cliffs, 3 edition, 1997.

[8] Pablo E Jojoa, Max Gerken, and Felipe M Pait. Um algoritmo acelerador de parâmetros. In XVI Simpósio Brasileiro de Telecomunicações, Vitória, —textsces, setembro 1999.

[9] Donald E Knuth. Selected papers in computer science. Cambridge University Press, 1996.

[10] Ilya Kolmanovsky and N Harris McClamroch. Developments in nonholonomic control problems. IEEE Control Systems Magazine, pages 20–36, December 1995.

[11] Iven Mareels and Jan Willem Polderman. Adaptive Systems: An Introduction. Birkhäuser, Boston, 1996.

[12] O Mayr. Zur Frügeschichte der Technischen Regelungen. Oldenbourg, München, 1969. (English translation: The origins of feedback control, MIT Press, Cambridge, Mass., 1970).

[13] A Stephen Morse and Felipe M Pait. mimo design models and internal regulators for cyclically-switched parameter-adaptive control systems. ieee Trans. Automatic Control, 39(9):1809–1818, September 1994.

[14] Felipe M Pait. Elementos de controle adaptativo. Universidade de São Paulo, 1998. (tese de livre-docência).

[15] Felipe M Pait. A tuner that accelerates parameters. Systems & Control Letters, 35(1):65–68, August 1998.

[16] Felipe M Pait and Fuad Kassab Jr. Parallel algorithms for adaptive control: Robust stability, volume Control Using Logic-Based Switching of Lecture Notes in Control and Information Sciences, 222, pages 262– 276. Springer Verlag, 1996.

[17] Felipe M Pait and Fuad Kassab Jr. On a class of switched, robustly stable, adaptive systems. In Conference on Decision and Control, Phoenix, Arizona, December 1999.

[18] Felipe M Pait and A Stephen Morse. Global tunability of one-dimensional siso systems. ieee Trans. Automatic Control, 37(10):1605–1608, October 1992.

[19] Felipe M Pait and A Stephen Morse. A cyclic switching strategy for parameter adaptive control. ieee Trans. Automatic Control, 39(6):1172– 1183, June 1994.

[20] Benedetto Piccoli, Felipe M Pait, and Anselmo Bittar. A hybrid controller for a nonholonomic system. Controle & Automação, 9(2):85–89, maio 1998.

[21] Ali Hussein Sayed, Thomas Kailath, and Babak Hassibi. Indefinite Quadratic Estimation and Control: A Unified Approach to H_{2} and H_{∞} Theories. SIAM, Philadelphia, 1999.